OpenStudy (anonymous):

What is the equation of the line of symmetry for the graph of y = –2(x – 4)^2 + 3? x = –2 x = 2 x = 3 x = 4

OpenStudy (anonymous):

I think its B

OpenStudy (jdoe0001):

can you change yourequation to look like => \(\large (x-h)^2=4p(y-k) \quad ?\)

OpenStudy (jdoe0001):

well... more like \(\large (x-h)^2=a(y-k)\)

OpenStudy (anonymous):

\[y=-2(x-4)^2+3\]

OpenStudy (jdoe0001):

hmm shouldn't matter actually... anyhow notice where the vertex is at http://nayak2010.files.wordpress.com/2011/01/picture-of-vertex-form-of-parabola.png

OpenStudy (anonymous):

yea

OpenStudy (jdoe0001):

so what do you think is the line of symmetry? do you know what line of symmetry means?